What is the volume of 3.5 x 10^24 molecules of CO2


1
Expert's answer
2021-03-29T06:01:08-0400

1 mole of CO2 has 6.02 x 10^23 molecules of CO2.

Say moles of CO2 has 3.5x10^24 molecules of CO2.

Therefore x = 3.5 ×10^24 /6.02 ×10^23 = 5.814

V=nm

r

V=nmr

= 5.814 × (12 + 2x16) = 5.814 × 44

= 255.816 ml